The Phonex Suns won two more thaqn three times as many games as they lost. The Suns played 82 games. How many wins and losses did the team have?
Answer by Earlsdon(6294)

Let W = the number of games won by the Phoenix Suns and L = the number of games lost.
W+L = 82 "The Suns played 82 games." Rewrite this as W = 82-L and substitute into the second equation.
W = 3L+2 "The Phoenix Suns won two more than three times as many games as they lost." Substitute W = 82-L.
82-L = 3L+2 Add L to both sides.
82 = 4L+2 Subtract 2 from both sides.
80 = 4L Divide both sides by 4.
20 = L and...
W = 3L+2
W = 3(20)+2
W = 62
The Phoenix Suns won 62 games and lost 20 games.